OpenStudy (anonymous):

using the quadratic formula solve each equation. give two forms of each solution an expressionwith radical and a calculator approximation rounded off to two decimal places. 2x^2+3x-4=0. I am a bit confused what the quadratic farmula is. Can anyone help.

OpenStudy (anonymous):

\[quadratic formula = -b \left( +or- \right)*(\sqrt{b ^{2}-4ac})/2a\]

OpenStudy (anonymous):

there is a song: negative b, plus or minus square root, b squared minus 4 a c, all over 2 a. kinda lame but good to memorize, just say each section twice ^^

OpenStudy (anonymous):

That is cool

OpenStudy (anonymous):

so how do I use this to solve my question?

OpenStudy (anonymous):

2x^2 +3x -4 = 0 a = 2 b = 3 c = -4 you just use the coefficients of the numbers

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I don't know how to set this up I am so lost